In pipe fitting you use the set,run,and travel of the right triangle to solve missing numbers,example: (A)square + (B)square = (C)square; well my problem is when we also use the numbers in a right triangle for squaring right triangles so if 3= the run and 4= the set and 5= the travel that's ok but when using set= travel x sine you come up with this, Travel x Sine = Set , so using 5 as the travel x
0.707,Sine of 45 =3.535 this is a have of inch more than the actual size so what is my problem

Answer by rothauserc(4718)

we have a 3, 4, 5 right triangle where side 4 is the set, side 3 is the run and side 5 is the travel. We need to find the acute angles in this triangle.
Let angle A be the angle opposite side 4(set) and angle B be opposite side 3(run). Then,
sin(-1) = 4/5 = 53.13 degrees = angle A
sin(-1) = 3/5 = 36.87 degrees = angle B
now
sin(A) = set / travel
set = travel * sin(53.13) = 5 * 0.799998928 = 3.99999464 approx 4